Masse of Li2CO3 necessary to prepare solution of 250,0ml with a 0,375mol/l concentration for Li+

1
Expert's answer
2021-01-19T04:39:26-0500

1) n / 0.250 = 0.375 / 1, n = 0.09375 - the amount of substance in 250 ml.


2) n = n / 2, n = 0.046875, because one molecule of the sault dissociates, creating 2 Li+ ions


3) n = m / M, m = nM = 0.046875 * 73.891 = 3.46 (grams).


SOLUTION: 3.46 grams.